Find the equation of the locus when the distance of P from the line x=4 is equal to its distance from the line y = 1
Answer by ikleyn(52799)

.
Find the equation of the locus when the distance of P from the line x=4 is equal to its distance from the line y = 1
-----------------------------------------------------------------
The locus for is the straight line passing through the point (x,y) = (4,1) and having the slope 1:

   y - 1 = x - 4,    or    y = x - 3.

Make a sketch. Draw the coordinate axes. Draw the line x = 4 and the line y = 1. Then draw your line.